OpenStudy (anonymous):

Can someone please help me to understand this? Aim: To determine the value of 'g' with the use of the simple pendulum. Thank you.

OpenStudy (anonymous):

' From the second law: \[mg \sin \theta=\frac{ mdv }{ dt }\]where the velocity v can be written in terms of \[v=\frac{ Ld \theta }{ DT }\]

OpenStudy (michele_laino):

@kym02 Please note that, formula that expresses period of little oscillations of a simple pendulum, is this: \[T=2 \pi \sqrt{\frac{ l }{ g }}\] wher g is, as usually, the gravity, namely g=981 cm/sec. Now squaring both sides of the above formula, we have: \[T ^{2}=4 \pi ^{2} \frac{ l }{ g }\] from which we can write: \[g=4 \pi ^{2}\frac{ l }{ T ^{2} }\], so you can measure g using measures of period T

OpenStudy (michele_laino):

@kym02 of course T is the period of little oscillations, and l is the length of your simple pendulum

OpenStudy (mrnood):

the above formula can be derived from the equations that @kym02 posted However - I seem to recall that it requires an assumption that sin theta = theta for small values of theta (in radians). The above formula is ONLY valid for small amplitude of swing for the pendulum

OpenStudy (anonymous):

Thank you very much. I now understand!😃